# Imperfect Competition in Markets for Short-Circuit Current Services

### v2

@@ Abstract
An important limitation of Inverter-Based Resources (IBR) is their reduced contribution to Short-Circuit Current (SCC), as compared to that of Synchronous Generators (SGs). With increasing penetration of IBR in most power systems, the reducing SCC poses challenges to a secure system operation, as line protections may not trip when required. In order to address this issue, the SCC ancillary service could be procured via an economic mechanism, aiming at securing adequate SCC on all buses. However, the suitability of markets for SCC services is not well understood, given that these could be prone to market-power issues: since the SCC contributions from various SGs to a certain bus are determined by the electrical topology of the grid, this is a highly local service. It is necessary to understand if SGs at advantageous electrical locations could exert market power and, if so, how it could be mitigated. In order to fill this gap, this paper adopts an SCC-constrained bilevel model to investigate strategic behaviors of SGs. To address the non-convexity due to unit commitment variables, the model is restructured through a primal-dual formulation. Based on a modified IEEE 30-bus system, cases with strategic SGs placed at different buses are analyzed. These studies demonstrate that agents exerting market power could achieve up to triple revenues from SCC provision, highlighting the need to carefully design these markets.
@@
    Index, Set of commitment decision pairs of SGs
@@ 1 Introduction
Eelectric power systems throughout the world are shifting to renewable energy dominated architectures for achieving the goal of net-zero carbon emissions in coming decades. This requires the massive installation of Inverter-Based Resources (IBR), mainly wind and solar generators.
@@
Market power issues in the electricity industry have been extensively analyzed. Authors in [9] investigated if market power is one of the reasons for the price increases in the wholesale electricity market of India. Other regions have also been examined, including California [10, 11], England and Wales [12], and the Nordic wholesale markets [13]. The relationship between market power, congestion, and renewable energy has also been explored, such as in the Italian day-ahead market [14]. Market power mitigation mechanisms in electricity markets have been proposed [15], including applying price caps as an efficient way to mitigate market power in terms of transmission congestion [16]. These efforts, among many others, demonstrate the importance of understanding if certain agents may exert market power in electricity systems.
Regarding modeling techniques for capturing imperfect competition effects, bilevel optimization has proven to be a powerful tool. Through a bilevel model one can simulate a situation where a leader makes decisions before the followers, who then react to those decisions [17]. Leaders, or price makers, are the agents with the capability to exert market power, due for example to their larger size in comparison to others, or their privileged location within a network. Mathematically, bilevel models are generally solved through replacing lower-level problems (representing the followers) by their equivalent Karush-Kuhn-Tucker (KKT) conditions, which are then included as constraints in upper level, therefore leading to an equivalent single-level model [17]. However, this approach requires the lower-level problem to be continuous and convex, as otherwise the KKT conditions either do not exactly correspond to the original problem, or simply can not be derived.
@@
In this work, we start by describing SCC constraints and verifying their effectiveness in a linear form. Then, for each network at hand, buses that face potential short-circuit risk are identified, and marginal prices for SCC services offered by various SGs are computed. Finally, the UC bilevel problem representing the market is solved via a primal-dual formulation, allowing to analyze the impact of strategic operation of different thermal units on the market outcome.
@@
    For the first time, a market formulation considering SCC as a service is proposed, by means of a bilevel optimization model. The non-convexity in the problem, due to the binary commitment state of SGs, is addressed via a primal-dual reformulation.
@@
    The proposed model allows to provide the first analysis of imperfect competition in markets for SCC services. We demonstrate that strategic SGs could achieve up to triple revenues from SCC markets, highlighting the need to carefully design these markets.
The remainder of this paper is structured as follows: Section 2 describes modeling of the SCC constraint and its application in the bilevel model. Section 3 explains how to tackle the non-convexity of the model by primal-dual reformulation. Section 4 includes relevant case studies to analyze market power issues, while Section 5 gives the conclusion and suggests lines for future work.
@@ 2 Short-Circuit Current Constrained Bilevel Problem Modeling
This section introduces the SCC constraints that are next applied into a bilevel model to secure all buses to an acceptable level of SCC through a market setting. Note that only three-phase nodal short-circuit faults are considered. For the sake of clarity, time index ‘$t$’ is omitted in the SCC expressions.
@@ 2.1 Formulation of Short-Circuit Current Constraints
The impedance matrix is calculated by inverting the corresponding admittance matrix, $Z_{ij}$ is thus hard to include in a duality-involved optimization problem. Therefore, a data-driven method is adopted to approximate the actual SCC constraints [8]. Coefficients $\{k_{b_{g}},k_{b_{c}},k_{b_{m}}\}$ are introduced to approximate the actual values of SCC (1) into:
@@
where (2a) is the approximate value of SCC, which is forced to be higher than $\textrm{I}_{b_{\textrm{lim}}}$. $\eta_{m}$ captures the interactions between pairs of SGs to simulate the nonlinearity, defined as (2b)-(2c). The approximate coefficients $k_{(\cdot)}$ are determined by operating an optimization-based classification procedure, comprising of a data generation process and a minimization problem that tries to reduce the error led by the approximation. This classification only serves to train the coefficients in (2) for fitting the actual SCC, meaning that it is in fact an offline process to the bilevel optimization model. An example for the training process used in this work can be accessed in [26].
@@ 2.2 Bilevel Model with Short-Circuit Current Constraints
The coefficients $\{k_{b_{g}},k_{b_{c}},k_{b_{m}}\}$ in (2a) are constants in the bilevel model as discussed above, therefore italics are no longer used for them. Besides, the number of dual variables corresponding to the equations for linearizing the quadratic term $\eta_{m}$ increases exponentially with more SGs involved, leading to high computational complexity. Therefore, $\eta_{m}$ is removed from the SCC constraint in the following models, still yielding an effective formulation as demonstrated in Section 4.2.
@@ 2.2.1 Upper-Level Model
Strategic SGs maximize the profits at the Upper-Level (UL) problem, where the optimal bidding decisions are determined to form the bids reported to the lower level. The UL model is expressed as:
@@
where objective function (3a) indicates that the revenues of strategic SGs come from the provision of energy and SCC, meanwhile the costs include no-load costs, marginal generation costs, start-up and shut-down costs. Formula (3b) shows the variables in UL problem. Constraints (3c)-(3d) confine the strategic multipliers of bids.
@@ 2.2.2 Lower-Level Model
where the system operating costs are represented in (4a), including the costs explained in (3a), the costs of SCC service and energy provided by IBR. (4b) includes primal variables in the LL model which follow constraints: SCC on bus $b$ contributed by SGs and IBR (4c); Demand-supply power balance constraints (4d); Minimum stable generation and maximum generation constraints for SGs (4e)-(4f); Ramp-up/down constraints (4g)-(4h); Start-up/shut-down costs (4i)-(4p); Minimum and maximum generation constraints for IBR (4q); (4r) impose integrality for commitment variables of SGs.
We compute the SCC service offer $\textrm{O}_{g,b,t}^{\textrm{SCC}}$ via a marginal unit pricing [7], from which the economic value of SCC from SG $g$ for bus $b$, i.e., $\textrm{k}_{bg}$, is obtained as the commitment price (this indicates that the service offer is implicitly expressed as $\textrm{O}_{g,b,t}^{\textrm{SCC}}/\textrm{k}_{bg}$, with the unit being €/p.u.). This computation implies obtaining the difference in objective values between two problems. First, problem (4) is solved as an energy-only UC without SCC-related terms in the objective. Second, the SCC provision $\textrm{k}_{bg}$ is eliminated in (4c), then calculating hourly operation costs of the modified dispatch problem. The difference in those two solutions represents the cost of forcing other units to startup for providing SCC. This whole process is further explained through a numerical example in Section 4.3.
@@ 3.1 Dualization of Primal Lower-Level Problem
where (6a) is the objective function of the dual problem. (6b) shows the dual variables involved. Constraints (6c)-(6f), (6g)-(6j), (6k)-(6n) and (6o) represent dual constraints regarding primal variables: $u_{\hat{g},t},u_{\check{g},t}$; $P_{\hat{g},t},P_{\check{g},t}$; $C_{\hat{g},t}^{\textrm{st}},C_{\hat{g},t}^{\textrm{sh}},C_{\check{g},t}^{\textrm{st}},C_{\check{g},t}^{\textrm{sh}}$ and $P_{c,t}$, respectively. Constraints (6p) express non-negativity of the dual variables. Note that $\lambda_{t}^{\textrm{E}}$ is also lower bounded by zero here to ensure non-negative energy revenues for SGs.
@@ 3.2 Primal-Dual Single-Level Optimization Model
where (7) is a single-level problem. Function (7a) couples the maximization of UL objective (3a) and minimization of DG between primal and dual LL problems. The DG is defined as (7b). Formulas (7c) and (7d) clarify the variables and constraints involved. The positive penalty factor W governs the trade-offs of objectives. An appropriate value for this parameter must be found for the problem at hand, as we discuss in Section 4.4, for which we define a ratio associated with primal and dual LL problems, to evaluate the error of solutions as follows:
@@
where a small value of $r^{\textrm{DG}}$ indicates a minor gap between the primal and dual problems, i.e., a more accurate solution.

### v3

@@ Abstract
An important limitation of Inverter-Based Resources (IBR) is their reduced contribution to Short-Circuit Current (SCC), as compared to that of Synchronous Generators (SGs). With increasing penetration of IBR in most power systems, the reducing SCC poses challenges to a secure system operation, as line protections may not trip when required. In order to address this issue, the SCC ancillary service could be procured via an economic mechanism, aiming at securing adequate SCC on all buses. However, the suitability of markets for SCC services is not well understood, given that these could be prone to market power issues: since the SCC contributions from various SGs to a certain bus are determined by the electrical topology of the grid, this is a highly local service. It is necessary to understand if SGs at advantageous electrical locations could exert market power and, if so, how it could be mitigated. In order to fill this gap, this paper, for the first time, adopts an SCC-constrained bilevel model to investigate strategic behaviors of SGs. To address the non-convexity due to unit commitment variables, the model is restructured through a primal-dual formulation. Based on a modified IEEE 30-bus system, cases with strategic SGs placed at different buses are analyzed. These studies demonstrate that strategic agents exerting market power by manipulating service prices and extending operating periods could achieve up to triple revenues from SCC provision, which reduces market efficiency and would increase the financial burden on consumers. These findings highlight the need for careful market design, for which potential measures to mitigate these market power issues are also discussed.
@@
    Index, Set of commitment decision pairs
@@
  • $\omega,\Omega$
    Index, Set of system operating points
@@ 1 Introduction
1.1 Background and Motivation
Electric power systems throughout the world are shifting to renewable energy dominated architectures for achieving the goal of net-zero carbon emissions in coming decades. This requires the massive installation of Inverter-Based Resources (IBR), mainly wind and solar generators.
@@ 1.1 Background and Motivation
Against this backdrop, this paper aims to explore, for the first time, two challenges in SCC service procurement: First, to examine potential strategic behaviors of SGs in SCC service markets; Second, to quantify the extent of market power exploitation by advantageous-location SGs. Analyzing such market power issues in SCC services could provide insights into the following concerns:
  1. 1.
    The markets for other stability services, such as frequency reserve and inertia, are already established in several real-world systems, such as Great Britain [9]. Thus, analyzing the potential market power risks in an SCC market is practically relevant today, as it could address a key need for SCC in modern power systems that can be met through market mechanisms.
  2. 2.
    If such risks exist, how they could be mitigated or avoided.
1.2 Review of Relevant Studies
Market power issues in the electricity industry have been extensively analyzed. Authors in [10] investigated if market power is one of the reasons for the price increases in the wholesale electricity market of India. Other regions have also been examined, including California [11, 12], England and Wales [13], and the Nordic wholesale markets [14]. The relationship between market power, congestion, and renewable energy has also been explored, such as in the Italian day-ahead market [15]. Market power mitigation mechanisms in electricity markets have been proposed [16], including applying price caps as an efficient way to mitigate market power in terms of transmission congestion [17]. These efforts, among many others, demonstrate the importance of understanding if certain agents may exert market power in electricity systems. However, the market for SCC services is not included in these studies, leaving unexamined whether and how SGs at advantageous electrical positions would exploit market power for their own interests.
Regarding modeling techniques for capturing imperfect competition effects, bilevel optimization has proven to be a powerful tool. Through a bilevel model one can simulate a situation where a leader makes decisions before the follower, who then reacts to those decisions [18]. Such a leader-follower structure, also known as a Mathematical Program with Equilibrium Constraints (MPEC), is well suited to capturing the interactions between strategic SGs and system operators in SCC service markets, providing a theoretical tool to fill the research gap described previously. Leaders, or price makers, are the agents with the capability to exert market power, due for example to their larger size in comparison to others, or their privileged location within a network. Mathematically, bilevel models are generally solved through replacing the lower-level problem (representing the follower) by its equivalent Karush-Kuhn-Tucker (KKT) conditions, which are then included as constraints in upper level, therefore leading to an equivalent single-level model [18]. However, this approach strictly requires the lower-level problem to be continuous and convex, as otherwise the KKT conditions either do not exactly correspond to the original problem, or simply can not be derived. Since SGs introduce discrete unit commitment variables, the traditional bilevel solution methods are not applicable in this work.
@@ 1.2 Review of Relevant Studies
1.3 Research Flow and Main Contributions
In this work, we start by describing SCC constraints and verifying their effectiveness in a linear form, laying the foundation for formulating the SCC market model. Then, we identify buses at risk of insufficient SCC and compute marginal prices for SCC services via the marginal unit pricing method proposed by [7], establishing a baseline to quantify perfect-competition market outcomes. Finally, the UC bilevel problem, which represents both energy and SCC service markets, is solved via a primal-dual formulation, enabling strategic analysis of how SGs at advantageous electrical positions exploit market power, and how such behavior impacts market clearing.
@@ 1.3 Research Flow and Main Contributions
    This paper presents the first market formulation of SCC as a service using a bilevel optimization model, where the non-convexity arising from the binary commitment of SGs is handled through a primal–dual reformulation.
@@
    The proposed model enables the first analysis of imperfect competition in SCC service markets, showing that strategic SGs can earn up to three times higher revenues by raising ancillary service prices and operating for longer periods, which underscores the importance of careful market design.
The remainder of this paper is structured as follows: Section 2 describes modeling of the SCC constraint and its application in the bilevel model. Section 3 explains how to tackle the non-convexity of the model by primal-dual reformulation. Section 4 includes relevant case studies to analyze market power issues and potential measures to mitigate such issues, while Section 5 gives the conclusion and suggests lines for future work.
@@ 2 Short-Circuit Current Constrained Bilevel Problem Modeling
This section introduces the SCC constraints that are next applied in a bilevel model to secure all buses to an acceptable level of SCC through a market setting. Note that only three-phase nodal short-circuit faults are considered. For the sake of clarity, time index ‘$t$’ is omitted in the SCC expressions.
@@ 2.1 Formulation of Short-Circuit Current Constraints
The impedance matrix is calculated by inverting the corresponding admittance matrix, $Z_{ij}$ is thus hard to include in a duality-involved optimization problem. Therefore, a data-driven method is adopted to approximate the actual SCC constraints [8]. Coefficients $\mathcal{K}=\{k_{b_{g}},k_{b_{c}},k_{b_{m}}\}$ are introduced to approximate the actual values of SCC (1) into:
@@
where (2a) is the approximate value of SCC, which is forced to be higher than $\textrm{I}_{b_{\textrm{lim}}}$. $\eta_{m}$ captures the interactions between pairs of SGs to simulate the nonlinearity, defined as (2b)-(2c). The approximate coefficients $\mathcal{K}$ are determined by operating an optimization-based classification procedure, comprising an enumeration for system operating points and a minimization problem that tries to reduce the error led by the approximation. This optimization is operated as follows:
$\min_{\mathcal{K}}\quad\sum_{\omega\in\Omega_{2}}\left(I_{b_{\textrm{L}}}^{(\omega)}-I_{b_{\textrm{SC}}}^{(\omega)}\right)^{2}$ (3a)
subject to:
$I_{b_{\textrm{L}}}^{(\omega)}<\textrm{I}_{b_{\textrm{lim}}}\quad\forall\,\omega\in\Omega_{1}$ (3b)
$I_{b_{\textrm{L}}}^{(\omega)}\geq\textrm{I}_{b_{\textrm{lim}}}\quad\forall\,\omega\in\Omega_{3}$ (3c)
$\Omega=\Omega_{1}\cup\Omega_{2}\cup\Omega_{3}$ (3d)
$\Omega_{1}=\left\{\omega\in\Omega\,\middle|\,I_{b_{\textrm{SC}}}^{(\omega)}<\textrm{I}_{b_{\textrm{lim}}}\right\}$ (3e)
$\Omega_{2}=\left\{\omega\in\Omega\,\middle|\,\textrm{I}_{b_{\textrm{lim}}}\leq I_{b_{\textrm{SC}}}^{(\omega)}<\textrm{I}_{b_{\textrm{lim}}}+\nu\right\}$ (3f)
$\Omega_{3}=\left\{\omega\in\Omega\,\middle|\,\textrm{I}_{b_{\textrm{lim}}}+\nu\leq I_{b_{\textrm{SC}}}^{(\omega)}\right\}$ (3g)
where (3a) is the objective for a minimum training error. $\Omega$ stands for all enumerated system operating points. Eqs. (3b) and (3e) enable the correct classification for all sub-limit SCC level. A positive parameter $\nu$ is introduced to define $\Omega_{2}$ and $\Omega_{3}$ (as specified in (3f)–(3g)), so that all data points in $\Omega_{3}$ are classified correctly and any misclassification is restricted to subset $\Omega_{2}$. Furthermore, $\nu$ should be set as small as possible while ensuring the feasibility, so that an approximation with desired training performance can be achieved, as analyzed in Section 4.2.
This classification only serves to train the coefficients in (2) for fitting the actual SCC, meaning that it is in fact an offline process to the bilevel optimization model. An example for the training process used in this work can be accessed in [27].
@@ 2.2 Bilevel Model with Short-Circuit Current Constraints
The coefficients $\mathcal{K}=\{k_{b_{g}},k_{b_{c}},k_{b_{m}}\}$ in (2a) are constants in the bilevel model as discussed above, therefore italics are no longer used for them. Besides, the number of dual variables corresponding to the auxiliary equations for linearizing the quadratic term $\eta_{m}$ increases exponentially with more SGs involved, leading to high computational complexity. Therefore, in the following models, we remove $\eta_{m}$ from the SCC constraint. Nevertheless, this simplifying assumption preserves the effectiveness of the formulation, as demonstrated in Section 4.2.
Note that the bilevel problem considered is an MPEC, that is, there is only one strategic player in the market who is able to manage multiple SGs, while all other participants are non-strategic players. Not simultaneously considering the strategic behaviors of all market participants allows us to capture the independent impact of this strategic player on the market and system operation, thereby decoupling the otherwise complex market problem. In fact, jointly considering multiple MPECs, corresponding to multiple strategic players participating in the market, would lead to an Equilibrium Problem with Equilibrium Constraints (EPEC). Interested readers are referred to [18].
@@ 2.2.1 Upper-Level Model
Strategic SGs maximize the profits from energy and SCC service markets at the Upper-Level (UL) problem, where the optimal bidding decisions are determined to form the bids reported to the lower level. The UL model is expressed as:
@@
where objective function (4a) indicates that the revenues of strategic SGs come from the provision of energy and SCC, meanwhile the costs include no-load costs, marginal generation costs, start-up and shut-down costs. Eq. (4b) shows the variables in UL problem. Eqs.  (4c)-(4d) confine the strategic multipliers of bids.
@@ 2.2.2 Lower-Level Model
where the system operating costs are represented in (5a), including the costs explained in (4a), the costs of SCC service and energy provided by IBR. Eq. (5b) includes primal variables in the LL model which follow constraints: SCC on bus $b$ contributed by SGs and IBR (5c); Demand-supply power balance constraints (5d); Minimum stable generation and maximum generation constraints for SGs (5e)-(5f); Ramp-up/down constraints (5g)-(5h); Start-up/shut-down costs (5i)-(5p); Minimum and maximum generation constraints for IBR (5q); Eq. (5r) imposes integrality for commitment variables of SGs.
We compute the SCC service offer $\textrm{O}_{g,b,t}^{\textrm{SCC}}$ via a marginal unit pricing [7], from which the economic value of SCC from SG $g$ for bus $b$, i.e., $\textrm{k}_{bg}$, is obtained as the commitment price (this indicates that the service offer is implicitly expressed as $\textrm{O}_{g,b,t}^{\textrm{SCC}}/\textrm{k}_{bg}$, with the unit being €/p.u.). This computation implies obtaining the difference in objective values between two problems. First, problem (5) is solved as an energy-only UC without SCC-related terms in the objective, denoted as $f^{*}(\textbf{x})$. Second, the SCC provision $\textrm{k}_{bg}$ is eliminated, as illustrated by (6), then calculating hourly operation costs of the modified dispatch problem, denoted as $f(\textbf{x})$. The difference in those two solutions (i.e., $f(\textbf{x})-f^{*}(\textbf{x})$) represents the cost of forcing other units to startup for compensating the removed SCC, i.e., the marginal service costs of SCC. This whole process is further explained through a numerical example in Section 4.3.
$\displaystyle\min\quad f(\textbf{x})$ (6a)
$\textrm{s.t.}~\sum_{g}\Delta\textrm{k}_{bg}u_{g,t}+\sum_{c}\textrm{k}_{bc}\upalpha_{c,t}\geq\textrm{I}_{b_{\textrm{lim}}}|\{\Delta\textrm{k}_{bg}|\textrm{k}_{bg},0\}$ (6b)
$\qquad g(\textbf{x})\leq 0$ (6c)
$\qquad h(\textbf{x})=0$ (6d)
where x represents decision variables. Eq. (6b) shows that the variable SCC coefficient $\Delta\textrm{k}_{bg}$ is removed by setting $\Delta\textrm{k}_{bg}=0$. Eqs. (6c)-(6d) indicate all other inequality and equality constraints remain unchanged.
@@ 3.1 Dualization of Primal Lower-Level Problem
where (8a) is the objective function of the dual problem. Eq. (8b) shows the dual variables involved. Eqs. (8c)-(8f), (8g)-(8j), (8k)-(8n) and (8o) represent dual constraints regarding primal variables: $u_{\hat{g},t}$$u_{\check{g},t}$; $P_{\hat{g},t}$, $P_{\check{g},t}$; $C_{\hat{g},t}^{\textrm{st}}$$C_{\hat{g},t}^{\textrm{sh}}$$C_{\check{g},t}^{\textrm{st}}$$C_{\check{g},t}^{\textrm{sh}}$ and $P_{c,t}$, respectively. Eq. (8p) expresses non-negativity of the dual variables. Note that $\lambda_{t}^{\textrm{E}}$ is also lower bounded by zero here to ensure non-negative energy revenues for SGs, otherwise they would not comply with the dispatch due to insufficient financial incentives. Meanwhile, it provides a reasonable lower bound for linearizing the bilinear term, as shown in the Appendix A.
@@ 3.2 Primal-Dual Single-Level Optimization Model
where (9) is a single-level problem. The objective function (9a) couples the maximization of UL objective (4a) and minimization of DG between primal and dual LL problems. The DG is defined as (9b). Eqs. (9c)-(9d) clarify the variables and constraints involved. The positive penalty factor W governs the trade-offs of objectives, i.e., the accuracy of market clearing and the consideration of profits. An appropriate value for this parameter must be found for the problem at hand, since a value too small for W would not sufficiently penalize the DG, while a value too large would instead drive the algorithm to overly focus on the DG minimization, and render the profit of strategic agents in (9) less important, potentially leading to less profitable strategic behaviors, as demonstrated in Section 4.4.1. Therefore, we define a ratio associated with primal and dual LL problems, to evaluate the error of solutions as follows:
@@
where a small value of $r^{\textrm{DG}}$ indicates a minor gap between the primal and dual problems, i.e., a more accurate solution of market clearing.